# The Serial Scaling Hypothesis

### v1

@@ Appendix A $\mathsf{TC}^{0}$ and $\mathsf{TC}$ Classes
Appendix B Computation capabilities of modern machine learning methods
@@
In this section, we point out an observational fact that machine learning methods that allow parallelization of their computation graph, such as MLPs, Transformers, and SSMs, have so far been shown to lack the ability to solve inherently serial problems. On the contrary, the only methods that are capable of solving serial problems are those that cannot be parallelized, such as RNNs, repeating layers, and Chain-of-Thought (CoT).
@@
In terms of computational complexity, multi-layer perceptrons (MLPs) have been formalized as a family of Boolean circuits with threshold gates called $\sf{TC}^{0}$ [85]. Consider a family of MLPs with constant $\mathcal{O}(1)$ depth, $\mathsf{poly}(n)$ neurons per layer, and $\mathcal{O}(1)$ numerical precision. A problem is in $\sf{TC}^{0}$ if and only if it is decidable by one forward pass through an MLP under this setup.
@@
It is widely suspected that $\sf{TC}^{1}$ is strictly larger than $\sf{TC}^{0}$. Therefore, any $\sf{TC}^{1}$-hard problem requires computations deepening at a rate of at least $\mathcal{O}(\ln n)$. We may consider the problem as being serial in a weaker, logarithmic sense. A simple example of $\sf{TC}^{1}$ is the word problem of the symmetric group on 5 elements $S_{5}$: Given $g_{1},g_{2},\dots,g_{n}\in S_{5}$, find $\prod_{i}g_{n}$ [63]. Intuitively, this is because $S_{5}$ is not a solvable group, and therefore there is no better way to multiply its elements than via a binary tree, which has $\mathcal{O}(\ln n)$ layers. Attempting to perform this with just $\mathcal{O}(1)$ layers would effectively require one to memorize the entire $S_{5}^{n}\to S_{n}$ multiplication table, which scales exponentially as $e^{n\ln(120)}$, an exponential depth-width trade-off.
@@
To solve the word problem of $S_{5}$ by a Transformer in one forward pass, we present it with input tokens $g_{1},g_{2},\dots,g_{n},y$, where $y$ is a special token. The Transformer’s output at $y$ is read out as the answer. Similarly, we may solve such a problem by presenting the same input tokens to a State Space Model (SSM) or a Recurrent Neural Network (RNN). Since a Transformer with $\mathcal{O}(1)$ layers and $\mathsf{poly}(n)$ dimensions can only solve problems in $\sf{TC}^{0}$ [71] in one forward pass, it cannot solve the word problem of $S_{5}$ that lies outside $\sf{TC}^{0}$.
@@
On the contrary, for an RNN, we can write the multiplication table of $S_{5}^{2}\to S_{5}$ directly into its weights, so it can solve the problem by unrolling for $\mathcal{O}(n)$ recurrence steps with $\mathcal{O}(1)$ layers and $\mathcal{O}(1)$ dimensions. Intuitively, the hidden states of an RNN keep track of the progress of multiplication as it performs the forward passes. However, RNN’s recurrence state dependency renders it non-parallelizable. Several families of SSMs were developed as a compromise that still have recurrence on hidden states, like an RNN, while making forward passes parallelizable, like a Transformer. The prototypical SSM architecture is Mamba [35], though there are many others.
@@
Unfortunately, it has been proven that there is not yet a “free serial-compute lunch,” in the sense that the main families of SSMs proposed so far still cannot solve the word problem of $S_{5}$ under the constraints of $\mathcal{O}(1)$ layers, $\mathsf{poly}(n)$ dimensions per layer, and one forward pass. Despite its apparent recurrence, the hidden state offered by an SSM is weaker in a computational sense than that offered by an RNN [110, 70, 74]. This theoretical fact rhymes with the empirical fact that in practice, MLPs, Transformers, and SSMs are all more parallelizable than RNNs.
@@
Only a genuinely serial method has been shown to go beyond the $\mathsf{TC}^{0}$ class. In addition to RNNs, this includes repeating layers and Chain-of-Thought (CoT). Repeating layers for $\mathcal{O}(\ln n)$ times enables a standard Transformer to solve tasks in the $\mathsf{TC}^{1}$ class [72]. With $\mathsf{poly}(n)$ CoT, requiring multiple forward passes before producing a final answer, a Transformer can be lifted from $\mathsf{TC}^{0}$ to $\mathsf{P}$ [31, 59]. As discussed in Section 4.5, the power of CoT has been well-attested in practice by the improved performance of reasoning models in complex math and science tasks.
@@
Appendix C Diffusion is in non-uniform $\mathsf{TC}^{0}$
In this section, we consider the non-uniform $\mathsf{TC}^{0}$ class, in contrast to the usual uniform classes. Note that uniform $\text{uniform }\mathsf{TC}^{i}\subseteq\text{non-uniform }\mathsf{TC}^{i}$. We prove that many diffusion models are restricted within that class. We need to assume non-uniformity because, at a certain point in the proof of the main theorem, we merely prove that something exists, without showing that it is also efficiently computable. We will highlight this in the proof.
C.1 Language modeling
@@
Comment. The theorem holds for any family of score-networks for which a single forward pass is in $\mathsf{TC}^{0}$. This includes, for example, Transformers and state-space models [71, 74].
@@
C.2 Main theorem
Theorem C.1.
@@ C.2 Main theorem
If this family solves a prefix language modeling problem at the limit of infinite time Score-Matching with Langevin Dynamics (SMLD) [107] with constant probability bound, then the problem is in the $\mathsf{TC}^{0}$ class.
@@
The idea of the proof is simple. We first construct a universal $\mathcal{O}(1)$-bound on how many steps are sufficient for sampling the SMLD within a constant probability bound, then we derandomize it according to [40], while still remaining within the $\mathsf{TC}^{0}$ class.
@@
By [57], there exists a constant $c>0$ that does not depend on any parameter in the statement of the theorem (a "universal constant"), such that there exists a noise schedule for DDPM that takes $T$ steps of the following kind. Using the schedule and the score-matching function $f_{\theta}$, we sample from a distribution $\rho_{DDPM,T}$. It satisfies the inequality
@@
Now, since $c$ is a universal constant, we can take $T$ to be $cc^{\prime}d^{2}$, to obtain an upper bound
@@
The key is that $T$ does not increase with $n$.
@@
Now we can derandomize this family, obtaining a $\mathsf{TC}^{0}$ family of Boolean circuits that solves the problem deterministically. The details of the derandomization method appear in [40, Prp. 4.2]. The big picture is as follows: we remove the probability by hard-coding a magic constant [^6] This is the part where we assume non-uniformity. In complexity theory, this is an “advice string”. per member of the family, such that sampling for polynomially many times, and taking the majority output, would always give the correct output. By Hoeffding’s inequality, such magic constants exist for a large enough polynomial.
@@
First, the full form of the inequality from [57] is
@@
Second, if we have no requirement on score-matching, then there is essentially no constraint on the computational power of SMLD, since any Turing machine can be constructed in SMLD with some (extremely non-score-matching) network. [^7] An explicit construction is published on Yuxi’s personal website. Yuxi would have published it on arXiv, but arXiv has put the paper “on hold” for months with no sign of actually publishing it.
@@
Appendix D Inherently Serial Problems in RL
@@ D.1 Definitions
Because a solution for $[\epsilon,b]$ yields one for $[\epsilon/b,1]$, it suffices to consider the special case $b=1$, which we call the $\epsilon$-approximate DO1 problem.
D.2 Non-parallelizability Results
We quote the following result from Kirousis and Spirakis [54].
Theorem D.1.
@@ D.2 Non-parallelizability Results
Theorem D.2 (optimal decision in the DO1 environment is inherently serial).
@@
[…]
@@
    We argue by contradiction. Assume there exists a parallel algorithm that, for some constant $10<\epsilon<$ 1, outputs a value function $V$ satisfying
[…]
@@
    Special case. We need to check first the special case where $d_{1}(C,x)=0$. This is done as follows. First, perform a topological sort of $C$, which is parallelizable (in $\mathsf{TC}^{2}$, in fact [21]). This then allows us to find all the gates that are in the first layer. Next, for each gate in the first layer, we test in parallel whether that gate outputs 1. Let their outputs be $y_{1},\dots,y_{m}$. If $\mathrm{OR}(y_{1},\dots,y_{m})=0$, then $d_{1}(C,x)=0$, because alternating circuits are monotonic. Otherwise, $d_{1}(C,x)\geq 1$.
@@
    Therefore the procedure is an approximation algorithm for DO1 that operates in parallel, contradicting Theorem D.1.
@@
    Assume for contradiction that there exists a parallel decision algorithm that is within an $\epsilon$ of optimality. That is, when the decision algorithm is applied in any DO1 environment, it always achieves reward at least $\epsilon r^{*}$. Then, for any particular DO1 environment, its first action must be an approximate solution as to whether $(C,x)$ or the chain possesses larger depth-of-1. Now, the same construction of (a) yields a parallel algorithm solving $\epsilon^{\prime}$-approximate DO1, again contradicting Theorem D.1.
@@
Although our analysis uses the DO1 problem in particular, there are some other $\mathsf{P}$-complete problems, such as linear programming, that are $\mathsf{P}$-complete to approximate. See for instance [91, 99, 25, 34] for some examples. This leads us to conjecture that inherent seriality may be a fairly common phenomenon in RL, not particular to the DO1 setup.
The barrier in Theorem D.2 can be bypassed in several ways:
@@
  • By employing a serial learning algorithm that runs in polynomial (not polylog) time, one may discover the right policy, which can then be compressed down into a policy that runs faster than serial. This bypasses the “$\mathsf{L}$-uniform” part of the obstacle.
  • By allowing the learned policy or value function to be non-parallel, the “$\mathsf{TC}$ circuit family” part of the obstacle is bypassed. This is true for certain RL algorithms, especially model-based methods that simulate the environment dynamics before making a decision.
@@
Part (b) of the theorem may explain the observation in Kevin et al. [53]. In that work, they trained both policy networks and value function networks by actor-critic methods, and noted that using deeper networks on both policy and value function improved performance. Part (b) of the theorem suggests that, when the network for approximating the value function contains less serial compute than the environment demands, then the corresponding exact $V$-maximizing policy would suffer arbitrarily bad worst-case performance. Though the theorem does not exactly apply when the policy is inexact, it suggests that the same phenomenon happens for an actual policy network trained to merely approximate the $V$-maximizing policy.

### v2

@@ Appendix D $\mathsf{TC}^{0}$ and $\mathsf{TC}$ Classes
Appendix E Computation capabilities of modern machine learning methods
@@
In this section, we point out an observational fact that machine learning methods that allow parallelization of their computation graph, such as MLPs, Transformers, and SSMs, have so far been shown to lack the ability to solve inherently serial problems. On the contrary, the only methods that are capable of solving serial problems are those that cannot be parallelized, such as RNNs, repeating layers, and CoT.
@@
In terms of computational complexity, MLPs have been formalized as a family of Boolean circuits with threshold gates called $\sf{TC}^{0}$ (Parberry & Schnitger, 1988). Consider a family of MLPs with constant $\mathcal{O}(1)$ depth, $\mathsf{poly}(n)$ neurons per layer, and $\mathcal{O}(1)$ numerical precision. A problem is in $\sf{TC}^{0}$ if and only if it is decidable by one forward pass through an MLP under this setup.
@@
It is widely suspected that $\sf{TC}^{1}$ is strictly larger than $\sf{TC}^{0}$. Therefore, any $\sf{TC}^{1}$-hard problem requires computations deepening at a rate of at least $\mathcal{O}(\ln n)$. We may consider the problem as being serial in a weaker, logarithmic sense. A simple example of $\sf{TC}^{1}$ is the word problem of the symmetric group on 5 elements $S_{5}$: Given $g_{1},g_{2},\dots,g_{n}\in S_{5}$, find $\prod_{i}g_{n}$ (Liu et al., 2022). Intuitively, this is because $S_{5}$ is not a solvable group, and therefore there is no better way to multiply its elements than via a binary tree, which has $\mathcal{O}(\ln n)$ layers. Attempting to perform this with just $\mathcal{O}(1)$ layers would effectively require one to memorize the entire $S_{5}^{n}\to S_{n}$ multiplication table, which scales exponentially as $e^{n\ln(120)}$, an exponential depth-width trade-off.
@@
To solve the word problem of $S_{5}$ by a Transformer in one forward pass, we present it with input tokens $g_{1},g_{2},\dots,g_{n},y$, where $y$ is a special token. The Transformer’s output at $y$ is read out as the answer. Similarly, we may solve such a problem by presenting the same input tokens to a State Space Model (SSM) or a Recurrent Neural Network (RNN). Since a Transformer with $\mathcal{O}(1)$ layers and $\mathsf{poly}(n)$ dimensions can only solve problems in $\sf{TC}^{0}$ (Merrill & Sabharwal, 2023b) in one forward pass, it cannot solve the word problem of $S_{5}$ that lies outside $\sf{TC}^{0}$.
@@
On the contrary, for an RNN, we can write the multiplication table of $S_{5}^{2}\to S_{5}$ directly into its weights, so it can solve the problem by unrolling for $\mathcal{O}(n)$ recurrence steps with $\mathcal{O}(1)$ layers and $\mathcal{O}(1)$ dimensions. Intuitively, the hidden states of an RNN keep track of the progress of multiplication as it performs the forward passes. However, RNN’s recurrence state dependency renders it non-parallelizable. Several families of SSMs were developed as a compromise that still have recurrence on hidden states, like an RNN, while making forward passes parallelizable, like a Transformer. The prototypical SSM architecture is Mamba (Gu & Dao, 2023), though there are many others.
@@
Unfortunately, it has been proven that there is not yet a “free serial-compute lunch,” in the sense that the main families of SSMs proposed so far still cannot solve the word problem of $S_{5}$ under the constraints of $\mathcal{O}(1)$ layers, $\mathsf{poly}(n)$ dimensions per layer, and one forward pass. Despite its apparent recurrence, the hidden state offered by an SSM is weaker in a computational sense than that offered by an RNN (Strobl et al., 2023; Merrill & Sabharwal, 2023a; Merrill et al., 2024). This theoretical fact rhymes with the empirical fact that in practice, MLPs, Transformers, and SSMs are all more parallelizable than RNNs.
@@
Only a genuinely serial method has been shown to go beyond the $\mathsf{TC}^{0}$ class. In addition to RNNs, this includes repeating layers and CoT. Repeating layers for $\mathcal{O}(\ln n)$ times enables a standard Transformer to solve tasks in the $\mathsf{TC}^{1}$ class (Merrill & Sabharwal, 2025). With $\mathsf{poly}(n)$ CoT, requiring multiple forward passes before producing a final answer, a Transformer can be lifted from $\mathsf{TC}^{0}$ to $\mathsf{P}$ (Feng et al., 2023; Li et al., 2024). As discussed in Section˜3.5, the power of CoT has been well-attested in practice by the improved performance of reasoning models in complex math and science tasks.
@@
Appendix F Diffusion is in non-uniform $\mathsf{TC}^{0}$
[…]
F.1 Language modeling
@@
Comment. The theorem holds for any family of score-networks for which a single forward pass is in $\mathsf{TC}^{0}$. This includes, for example, Transformers and state-space models (Merrill & Sabharwal, 2023b; Merrill et al., 2024).
@@
F.2 Main theorem
Theorem F.1.
@@ F.2 Main theorem
If this family solves a prefix language modeling problem at the limit of infinite time Score-Matching with Langevin Dynamics (SMLD) (Song & Ermon, 2019) with constant probability bound, then the problem is in the $\mathsf{TC}^{0}$ class.
Before we begin the proof, we give the intuitive sketch of the proof. It has two parts. The first part uses the fast convergence rate of the Euler–Maruyama method for solving the diffusion equation to show that we do not need many steps to correctly sample from the diffusion model with high probability.
$x\mapsto\text{majority}(f(x,s_{1}),f(x,s_{2}),\dots,f(x,s_{k(n)})).$
If $s$ is randomly sampled, then let the probability that $f(x,s)$ is wrong be upper-bounded by a constant $p<1/2$. By Hoeffding’s inequality (Hoeffding, 1963), the probability that the majority vote is correct is $\geq 1-e^{-2k(n)(p-1/2)^{2}}$.
Sample $k(n)$ random seeds, and fix them. This provides a deterministic model. Now, we try this deterministic model on every single possible input of length $n$. There are only $\text{vocab}^{n}$ of them. If we set $k(n)\geq\frac{\ln(\#\text{vocab})}{2(p-1/2)^{2}}$, then by the union bound, the probability that the majority vote is correct on all inputs of length $n$ is nonzero. Thus, there exists a specific choice of random seeds $(s_{1},\dots,s_{k(n)})$ that makes the compound model correct on all inputs of length $n$.
@@
The idea of the proof is simple. We first construct a universal $\mathcal{O}(1)$-bound on how many steps are sufficient for sampling the SMLD within a constant probability bound, then we derandomize it according to (Hajnal et al., 1993), while still remaining within the $\mathsf{TC}^{0}$ class.
@@
By (Li & Yan, 2024), there exists a constant $c>0$ that does not depend on any parameter in the statement of the theorem (a "universal constant"), such that there exists a noise schedule for DDPM that takes $T$ steps of the following kind. Using the schedule and the score-matching function $f_{\theta}$, we sample from a distribution $\rho_{DDPM,T}$. It satisfies the inequality
@@
Now, since $c$ is a universal constant, we can make up another sufficiently large universal constant $c^{\prime}$, and set $T$ to be $cc^{\prime}d^{2}$, to obtain an upper bound
@@
The key is that $T$ increases with $d$, but does not increase with $n$.
@@
Now we can derandomize this family, obtaining a $\mathsf{TC}^{0}$ family of Boolean circuits that solves the problem deterministically. The details of the derandomization method appear in (Hajnal et al., 1993, Prp. 4.2). The big picture is as follows: we remove the probability by hard-coding a magic constant [^6] This is the part where we assume non-uniformity. In complexity theory, this is an “advice string”. per member of the family, such that sampling for polynomially many times, and taking the majority output, would always give the correct output. By Hoeffding’s inequality, such magic constants exist for a large enough polynomial.
@@
First, the full form of the inequality from (Li & Yan, 2024) is
@@
Second, if we have no requirement on score-matching, then there is essentially no constraint on the computational power of SMLD, since any Turing machine can be constructed in SMLD with some (extremely non-score-matching) network. The explicit construction is published in (Liu, 2025).
@@
Appendix G Inherently Serial Problems in RL
@@ G.1 Definitions
Because a solution for $\epsilon,b]$ yields one for $\epsilon/b,1]$, it suffices to consider the special case $b=1$, which we call the $\epsilon$-approximate DO1 problem.
G.2 Non-parallelizability Results
We quote the following result from Kirousis and Spirakis (Kirousis & Spirakis, 1988).
Theorem G.1.
@@ G.2 Non-parallelizability Results
Theorem G.2 (optimal decision in the DO1 environment is inherently serial).
@@
    No parallel algorithm can compute an approximate optimal value function $V$ such that  $\exists\,0<\epsilon<b$ with $V(s)\in[\epsilon V^{*}(s),\,bV^{*}(s)]$ for every state $s$ in every DO1 environment.
@@
    We argue by contradiction. Assume there exists a parallel algorithm that, for some constant $0<\epsilon<1$, outputs a value function $V$ satisfying
    $V(s)\;\in\;[\epsilon\,V^{*}(s),\,V^{*}(s)]\quad\text{for every state }s\text{ in every DO1 environment}.$
@@
    Special case. We need to check first the special case where $d_{1}(C,x)=0$. This is done as follows. First, perform a topological sort of $C$, which is parallelizable (in $\mathsf{TC}^{2}$, in fact (Cook, 1985)). This then allows us to find all the gates that are in the first layer. Next, for each gate in the first layer, we test in parallel whether that gate outputs 1. Let their outputs be $y_{1},\dots,y_{m}$. If $\mathrm{OR}(y_{1},\dots,y_{m})=0$, then $d_{1}(C,x)=0$, because alternating circuits are monotonic. Otherwise, $d_{1}(C,x)\geq 1$.
@@
    Therefore the procedure is an approximation algorithm for DO1 that operates in parallel, contradicting Theorem G.1.
@@
    Assume for contradiction that there exists a parallel decision algorithm that is within an $\epsilon$ of optimality. That is, when the decision algorithm is applied in any DO1 environment, it always achieves reward at least $\epsilon r^{*}$. Then, for any particular DO1 environment, its first action must be an approximate solution as to whether $(C,x)$ or the chain possesses larger depth-of-1. Now, the same construction of (a) yields a parallel algorithm solving $\epsilon^{\prime}$-approximate DO1, again contradicting Theorem G.1.
@@
Although our analysis uses the DO1 problem in particular, there are some other $\mathsf{P}$-complete problems, such as linear programming, that are $\mathsf{P}$-complete to approximate. See for instance (Sahni & Gonzalez, 1976; Serna, 1991; Díaz et al., 1997; Greenlaw et al., 1995) for some examples. This leads us to conjecture that inherent seriality may be a fairly common phenomenon in RL, not particular to the DO1 setup.
The barrier in Theorem G.2 can be bypassed in several ways:
@@
  • By employing a serial learning algorithm that runs in polynomial (not polylog) time, one may discover the right policy, which can then be compressed down into a policy that runs faster than serial. This bypasses the $\mathsf{L}$-uniform” part of the obstacle.
  • By allowing the learned policy or value function to be non-parallel, the $\mathsf{TC}$ circuit family” part of the obstacle is bypassed. This is true for certain RL algorithms, especially model-based methods that simulate the environment dynamics before making a decision.
@@
Part (b) of the theorem may explain the observation in Kevin et al. (2025). In that work, they trained both policy networks and value function networks by actor-critic methods, and noted that using deeper networks on both policy and value function improved performance. Part (b) of the theorem suggests that, when the network for approximating the value function contains less serial compute than the environment demands, then the corresponding exact $V$-maximizing policy would suffer arbitrarily bad worst-case performance. Though the theorem does not exactly apply when the policy is inexact, it suggests that the same phenomenon happens for an actual policy network trained to merely approximate the $V$-maximizing policy.